Circular No. 5786

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1915+105 J. Greiner, Max-Planck-Institut für Extraterrestrische Physik, Garch-

ing, communicates: “Using a ROSAT HRI observation of the 1992 transient GRS 1915+105 in Aquila, its x-ray position has been determined to ∼ ± 10′′ accuracy: α = 19^{h}15^{m}11^{s}.8, δ = +10^{o}56′47′′ (equinox 2000.0). A check of the second Palomar Sky Survey plates (courtesy of N. Reid, Palomar Observatory) has revealed three highly reddened stars fainter than magnitude R = 21 within the 10′′ x-ray error box. Two of these objects are variable between 1992 June 10 and Aug. 8: one has brightened (as the x-ray source; cf. IAUC 5619), while the other has become fainter. The radio position (IAUC 5773) is ∼ 5′′ from the central position given above and coincides in position with the third, non-variable optical object. Optical spectroscopy is in progress.”

PSR 0540–693 A. Shearer, M. Redfern, T. Rowold, P. O’Kane, R. Butler, and C.

O’Byrne, University College, Galway; and H. Pedersen, Copenhagen University Observatory, communicate: “We have observed the position of the optical pulsar 0540–693 (Middleditch and Pennypacker 1985, Nature, 313,

659) using the European Southern Observatory’s New Technology Telescope (+ UCG/DIAS/ESO MAMA image sharpening camera, TRIFFID) on 1993 Jan. 15-16. On each night, the 19.838-Hz barycentric power spectrum component (computed from the pulsar period given in Manchester et al. (1993, Ap.J. 403, L103) showed a single spatial peak, consistent in height with Middleditch and Pennypacker, in the center of the nebulosity and coincident with the northern maximum of Caraveo et al. (1992, Ap.J. 395, L103).”

α PISCIS AUSTRINI V. Mannings and J. P. Emerson, Department of Physics, Queen Mary

and Westfield College, London, report: “We find a 3σ upper limit of 24 mJy at 1.3 mm for α PsA (Fomalhaut), the observation acquired in excellent conditions on Apr. 28.72-28.75 UT using the James Clerk Maxwell Telescope (+ UKT14) at Mauna Kea. A 19′′.8 beam and 90′′ chop in azimuth were used. This upper limit is a factor of 2 below the 1.3-mm detection reported on IAUC 5732.”

1993 May 10 Daniel W. E. Green